# Assisting Mathematical Formalization with A Learning-based Premise Retriever

### v2

@@ 3.3 Learning Algorithm
Based on the Masked Language Modeling (MLM) [36, 8] method, we pre-train both of the CFR and CAR modules on the formalized corpus we collected. Their tokenizers are the same and trained on formalized language corpus using WordPiece [35] algorithm. Then, we use the state-premises pairs to fine-tune the pre-trained modules separately by contrastive learning.
3.3.1 Contrastive Learning of CFR Module
@@
For $(\bm{s}_{i},\mathcal{P}_{i})\in\mathcal{D}$, we use $\bm{p}_{ij}\in\mathcal{P}_{i}$ along with negatives sampled from $\mathbb{M}$ to construct a set $\mathcal{P}_{ij}^{\prime}$, which contains one positive $\bm{p}_{ij}$ and $|\mathcal{P}_{ij}^{\prime}|-1$ negatives. The optimization objective of contrastive learning is formulated as
$\min\sideset{}{{}_{i=1}^{|\mathcal{D}|}}{\sum}\sideset{}{{}_{j=1}^{|\mathcal{P}_{i}|}}{\sum}-\log\frac{\exp(\text{sim}(\bm{s}_{i},\bm{p}_{ij})/\tau)}{\sum_{\bm{p}^{\prime}\in\mathcal{P}_{ij}^{\prime}}\exp(\text{sim}(\bm{s}_{i},\bm{p}^{\prime})/\tau)},$ (6)
@@
where $\tau$ refers to the temperature. We use the Homogeneous In-Batch Negative Sampling, which calls for many negative samples to guarantee the embedding’s discriminativenss [15, 37, 28]. In our work, this is implemented by the usage of in-batch negatives — for each query, the negatives are randomly sampled from the corpus, excluding the other positive premises of the query. Given a batch of $B$ samples, it results in $B\times|\mathcal{P}_{ij}^{\prime}|-1$ negative samples.
3.3.2 Contrastive Learning of CAR module
@@
The CAR module is learned in the same contrastive learning framework, in which the choice of negatives is crucial. We sample hard negatives from the top-$k_{1}$ premises selected by the retrieval model and they will be used in the re-ranking model training process. During testing, we also use the retrieval model to retrieve the top-$k_{2}$ premises and re-ranking them by the re-ranking model. The loss function used for training the re-ranking model is the cross-entropy loss, and the optimization objective is formulated as
$\min\sideset{}{{}_{i=1}^{|\mathcal{D}|}}{\sum}\sideset{}{{}_{j=1}^{|\mathcal{P}_{i}|}}{\sum}-\log\frac{Pr(\bm{s}_{i},\bm{p}_{ij})}{\sum_{\bm{p}^{\prime}\in\mathcal{P}_{ij}^{\prime}}Pr(\bm{s}_{i},\bm{p}^{\prime})},$ (7)
@@
where $\mathcal{P}_{ij}^{\prime}$ is made up by one positive $\bm{p}_{ij}$ and several hard negatives. $Pr(\bm{s},\bm{p})$ is the relevant probability obtained via Eq. (5).
3.3.3 Training a Tactic Generator for Evaluation
@@
After training the retrieval model, we integrate it with the generator for testing. Most prior work on generators [39, 34, 41] utilizes state-tactic pairs to fine-tune a pre-trained model for tactic generation. Leandojo [41] incorporates premise retrieval by leveraging the retrieval model to extract relevant premises based on the state from the training dataset. The premises are then concatenated with the state to form premise-augmented state-tactic pairs used for fine-tuning. However, this approach results in a high degree of coupling between the retrieval model and the generator, as prior knowledge of the retrieval result distribution is incorporated during training.
@@
To ensure a fair experimental setup, we aim to decouple the generator from the retrieval model and train the generator independently. As a substitution of the retrieval model, we randomly select the positive and negative premises from the library and prepend them to the state, thereby creating premise-augmented state-tactic pairs. For each tactic, we add up to ten premises. The number of positive premises selected is randomly determined, ranging from one to the lesser of the total available positive premises or ten. The remaining premises are negative. Half of these negatives are selected from the same module as the positive samples, while the remaining are chosen from other modules. Following [41], we use these pairs to fine-tune the ByT5 [40] model for tactic generation. The loss function we use is typically cross-entropy loss.
@@ 4 Experiments
To demonstrate the effectiveness of our retrieval method, we provide conducted experiments to answer the following research questions (RQs).
  • RQ1: How does our proposed retrieval model compare to state-of-the-art retrieval models in formal premise retrieval?
  • RQ2: How do our model architecture and training strategy affect the model performance?
  • RQ3: What is the robustness of our method with respect to variations in data and model hyperparameters?
  • RQ4: What is the impact of our retrieval model on the performance of retrieval-augmented theorem provers?
@@ 4.1 Experiment Setup
4.1.1 Implementation Details
Our retrieval and re-ranking models are based on the BERT architecture, which consists of 6 layers, each with 12 attention heads. It features a hidden size of 768 and an intermediate size of 3,072, complemented by a vocabulary size of 30,522. For the retrieval model, the maximum position embeddings are configured at 512. The maximum length for states is set at 512 and for premises’ arguments and goals set at 256 respectively. The batch size is set at 32 and $|\mathcal{P}_{ij}^{\prime}|$ mentioned in 3.3.1 is set at 2. In contrast, the re-ranking model has its maximum position embeddings set at 1,024. We set the batch size at 2, gradient accumulation steps at 8, and $|\mathcal{P}_{ij}^{\prime}|$ at 8. We construct the pre-training corpus by concatenating states from the training set and all premises’ argument lists and goals from the corpus. The results of our method presented in Table 3 are obtained by first training the retrieval model and selecting the top-$100$ results as hard negative candidates. We then train the re-ranking model and use it to reorder the top-$20$ results from the retrieval stage. The experiments are conducted on 8 RTX 4080 servers.
4.1.2 Baselines
As mentioned above, ReProver [41] is a model specifically trained on the Lean dataset, using formal states and premises. Therefore, it serves as our primary baseline. Moreover, we retrain the model using our dataset and following the setting in [41]. In addition to ReProver, we compare against several other commonly used retrieval models that have demonstrated strong performance in their specific domains.
  • UniXcoder-base [11]: It is one of the state-of-the-art code embedding models to transform code snippets into hidden vectors. It leverages multimodal data, such as code comments and abstract syntax trees (ASTs), to pre-train code representations.
  • E5-large-v2 [37]: It is trained in a contrastive manner using weak supervision signals derived from a curated, large-scale text-pair dataset (referred to as CCPairs) and demonstrates strong performance on several English retrieval tasks.
  • BGE-m3 [4]: It is distinguished for its versatility and excels in multiple functionalities, multilingual capabilities, and fine granularity retrieval tasks.
We fine-tune the three aforementioned baselines using our dataset. Following the parameters provided in [4], we set the learning rate to 1e-5 and use a linear scheduler. Due to server constraints, the batch sizes for the three models are set to 16, 8, and 6, respectively.
4.1.3 Metrics
To evaluate the performance of various retrieval models, we utilize four widely adopted metrics: Precision, Recall, F1-Score, and nDCG. Since the nDCG considers the retrieved results’ relevance and position, it allows for multiple relevance levels. We define the relevance criteria as shown in Table 1, which provides the relevance scores required for calculating nDCG.
@@ 4.1.3 Metrics
Rating Score Description
Match 1 Exact match to one of the premises used in the tactic.
@@
4.1.4 Data Split
To evaluate the performance of the model across different feature datasets, we employ four distinct data split strategies. In particular, we represent a proof as $P=\{t_{i}\}_{i=1}^{|P|}$, which is a sequence of tactics, and apply Proof Length and Premise Frequency to characterize the proof, i.e.,
@@ 4.1.4 Data Split
For each dataset, the training set contains 65,567 theorems, while the validation and test sets each consist of 2,000 theorems. The average proof length and premise frequency for each dataset are shown in Table 2.
Table 2: The average proof length and premise frequency of four different split strategies.
Split Proof Length Premise Frequency
train valid test train valid test
RD 2.48 2.54 2.44 2.25 2.31 2.15
RI 2.41 3.54 3.62 2.20 3.03 3.16
PL 2.35 4.50 4.77 2.26 2.12 2.11
PF 2.49 2.31 2.31 2.13 4.08 4.63
@@
Table 4: Ablation studies on pre-training, tokenizer, and similarity calculation on Random split dataset for the CFR module.
@@
Pre-train Tokenizer Similarity R@1 R@5 R@10 P@1 P@5 P@10 F@1 F@5 F@10 n@1 n@5 n@10
$\times$ $\times$ fine-grained in Eq. (3) 4.63 20.09 29.05 8.35 7.45 5.62 5.96 10.87 9.42 0.2443 0.3503 0.4030
$\times$ $\checkmark$ fine-grained in Eq. (3) 4.93 20.49 30.33 8.04 7.50 5.79 6.11 10.98 9.72 0.2368 0.3493 0.4034
$\checkmark$ $\times$ fine-grained in Eq. (3) 6.31 23.81 34.21 10.34 8.77 6.59 7.84 12.82 11.05 0.2484 0.3655 0.4240
$\checkmark$ $\checkmark$ conventional in Eq. (2) 5.65 25.87 36.78 9.36 9.55 7.12 7.05 13.95 11.92 0.2382 0.3662 0.4238
$\checkmark$ $\checkmark$ fine-grained in Eq. (3) 5.94 25.81 37.70 9.80 9.53 7.24 7.40 13.92 12.15 0.2422 0.3674 0.4293
4.2 Main Results (RQ1)
After training different models on the training dataset and evaluating them on the corresponding test sets, we obtain the results presented in Table 3. From this table, we can observe that our method demonstrates exceptional performance, significantly outperforming other baseline models. Specifically, UniXcoder-base [11], E5-large-v2 [15], and BGE-m3 [4] exhibit generally poor performance, which can be attributed to the substantial discrepancy between the pre-training corpora used by these models (based on natural language data) and the Lean language. Although these models are fine-tuned, they cannot fully capture the core aspects of the Lean language due to Lean’s unique syntax and semantic features and thus suffer suboptimal performance. As a result, these models likely require more specialized fine-tuning strategies and parameters tailored to the Lean language to improve their performance.
In contrast, the ReProver [41] model performs relatively well, likely due to its specifically designed parameter-tuning strategy. Furthermore, ReProver utilizes ByT5 [40], a byte-level tokenizer, which may allow the model to better handle the special symbols and structures in the Lean language. Our approach, on the other hand, improves performance by retraining the tokenizer on the Lean corpus and pre-training the retrieval model, allowing it to better adapt to the syntax and semantic features of Lean, resulting in a significant boost in overall performance.
@@ 4.2 Main Results (RQ1)
4.3 Ablation Studies (RQ2)
4.3.1 Impacts of Pre-Training, Tokenizer, and Similarity Calculation
We evaluate the impacts of pre-training the model, training the specific tokenizer for Lean, and the proposed similarity calculation in Table 4, respectively. To ensure a fair comparison, we filter out the impact of the re-ranking module and only compare the results provided by the retrieval module. This table shows a significant performance improvement when the model is pre-trained on the Lean corpus compared to not performing pre-training. This indicates that the pre-training process helps the model acquire semantic information about Lean in advance. The experiment also shows that if we only pre-train the model without training a new tokenizer, the performance when k=1 is slightly better but degrades a lot when k=5 or 10. This is because Lean has a formal language system, which differs from natural languages, highlighting the necessity of training a dedicated tokenizer for Lean. In addition, we evaluate the impacts of different similarity calculation methods on the retrieval results. The results in Table 4 emphasize the benefits of assessing similarity in a fine-grained manner.
@@ 4.3.1 Impacts of Pre-Training, Tokenizer, and Similarity Calculation
Figure 4: (a) The Recall@k of our method with or without re-ranking, compared with ReProver. (b) The performance and efficiency comparison on Random split datasets. The size of each point reflects the parameter size of each model.
4.3.2 The Effect of Re-ranking Model
@@
We validate the effectiveness of using the proposed re-ranking module. As shown in Figure 4 (a), when only the CFR module is used, our model outperforms the ReProver only for larger values of k in Recall@k (e.g., k $\geq 10$). Applying the CAR module improves the recall of our model when k $<10$, so that our model can consistently outperform ReProver and users are likely to find useful premises in a relatively short list. In other words, re-ranking contributes to a significant improvement in performance. Figure 4 (b) presents a comparison for our model and other baselines in terms of model size and inference GFLOPs. Without re-ranking, our model achieves the lowest GFLOPs and it outperforms in Recall@1 most of the baselines except ReProver. When re-ranking the top-$5$ retrieval results by our CAR module, our model outperforms all the baselines and its GFLOPs and model size are comparable to those of ReProver.
In principle, the CAR module helps improve the Recall of our model, but its complexity increases linearly with the number of results to re-rank. Therefore, the CAR module of our model is optional in practice, as there exists a trade-off between effectiveness and efficiency. The decision to re-ranking can be made based on factors such as sensitivity to the order of retrieval results or limitations on the information window. According to experimental results, when the number of retrieved results exceeds 50, the performance improvements obtained by re-ranking are limited, making it unnecessary. For user groups that are insensitive to the order of retrieval results, such as human theorem provers, it may be unnecessary to allocate additional computational resources for re-ranking. However, in the case of retrieval modules used in automated theorem provers, where high retrieval precision is required and computational cost is not a bottleneck, incorporating a re-ranking process can be justified. Note that, as a general retrieval technique, we can also apply the re-ranking procedure to other baselines. However, with larger backbone models (e.g., ReProver), they will need more computation, making it difficult to deploy for our search engine. How to further improve model performance and efficiency simultaneously is left as our future work.
Figure 5: The variation of the normalized values of the model’s metrics under different data perturbations.
4.4 Robustness Experiments (RQ3)
4.4.1 Effect of Data Perturbation

### v3

@@ 3.3 Learning Algorithm
Based on the Masked Language Modeling (MLM) (Taylor, 1953; Devlin et al., 2019) method, we pre-train the CFR and CAR modules on the formalized corpus we collected. Their tokenizers are the same and trained on formalized language corpus using WordPiece (Song et al., 2021) algorithm. Then, we use the state-premises pairs to fine-tune the pre-trained modules separately by contrastive learning.
@@
Contrastive Learning of CFR Module For $(\bm{s}_{i},\mathcal{P}_{i})\in\mathcal{D}$, we use $\bm{p}_{ij}\in\mathcal{P}_{i}$ along with negatives sampled from $\mathbb{M}$ to construct a set $\mathcal{P}_{ij}^{\prime}$, which contains one positive $\bm{p}_{ij}$ and $|\mathcal{P}_{ij}^{\prime}|-1$ negatives. The optimization objective of contrastive learning is formulated as
$\min\sum_{i=1}^{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{P}_{i}|}-\log\frac{\exp(\text{sim}(\bm{s}_{i},\bm{p}_{ij})/\tau)}{\sum_{\bm{p}^{\prime}\in\mathcal{P}_{ij}^{\prime}}\exp(\text{sim}(\bm{s}_{i},\bm{p}^{\prime})/\tau)},$ (5)
@@
where $\tau$ refers to the temperature. We use the Homogeneous In-Batch Negative Sampling, which calls for many negative samples to guarantee the embedding’s discriminativenss (Izacard et al., 2022; Wang et al., 2022; Qu et al., 2021). In our work, this is implemented by the usage of in-batch negatives — for each query, the negatives are randomly sampled from the corpus, excluding the other positive premises of the query. Given a batch of $B$ samples, it results in $B\times|\mathcal{P}_{ij}^{\prime}|-1$ negative samples.
@@
Contrastive Learning of CAR module The CAR module is learned in the same contrastive learning framework, in which the choice of negatives is crucial. We sample hard negatives from the top-$k_{1}$ premises selected by the retrieval model and they will be used in the re-ranking model training process. During testing, we also use the retrieval model to retrieve the top-$k_{1}$ premises and re-ranking them by the re-ranking model. The loss function used for training the re-ranking model is the cross-entropy loss, and the optimization objective is formulated as
$\min\sum_{i=1}^{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{P}_{i}|}-\log\frac{Pr(\bm{s}_{i},\bm{p}_{ij})}{\sum_{\bm{p}^{\prime}\in\mathcal{P}_{ij}^{\prime}}Pr(\bm{s}_{i},\bm{p}^{\prime})},$ (6)
@@
where $\mathcal{P}_{ij}^{\prime}$ contains one positive $\bm{p}_{ij}$ and some hard negatives. $Pr(\bm{s},\bm{p})$ is the relevant probability obtained via Eq. (4).
@@
Training Tactic Generator After training the retrieval model, we integrate it with the generator for testing. The generator is trained in an independent paradigm for a fair comparison between retrievers. Most prior work on generators (Welleck & Saha, 2023; Song et al., 2023; Yang et al., 2024) utilizes state-tactic pairs to fine-tune a pre-trained model for tactic generation. Leandojo (Yang et al., 2024) incorporates premise retrieval by leveraging the retrieval model to extract relevant premises based on the state from the training dataset. The premises are then concatenated with the state to form premise-augmented state-tactic pairs used for fine-tuning. However, this approach results in a high degree of coupling between the retrieval model and the generator, as prior knowledge of the retrieval result distribution is incorporated during training.
@@
To ensure a fair experimental setup, we aim to decouple the generator from the retrieval model and train the generator independently. As a substitution of the retrieval model, we randomly select the positive and negative premises from the library and prepend them to the state, thereby creating premise-augmented state-tactic pairs. For each tactic, we add up to ten premises. The number of positive premises selected is randomly determined, ranging from one to the lesser of the total available positive premises or ten. The remaining premises are negative. Half of these negatives are selected from the same module as the positive samples, while the remaining are chosen from other modules. Following (Yang et al., 2024), we use these pairs to fine-tune the ByT5 (Xue et al., 2022) model for tactic generation. The loss function we use is typically cross-entropy loss.
@@ 4.1 Experiment Setup
Implementation Details Our retrieval and re-ranking models are based on the BERT architecture, which consists of 6 layers, each with 12 attention heads. It features a hidden size of 768 and an intermediate size of 3,072, complemented by a vocabulary size of 30,522. For the retrieval model, the maximum position embeddings are configured at 512. The maximum length for states is set at 512 and for premises’ arguments and goals set at 256 respectively. The batch size is set at 32, and $|\mathcal{P}_{ij}^{\prime}|$ mentioned in 3.3 is set at 2. In contrast, the re-ranking model has its maximum position embeddings set at 1,024. We set the batch size at 2, gradient accumulation steps at 8, and $|\mathcal{P}_{ij}^{\prime}|$ at 8. We construct the pre-training corpus by concatenating states from the training set and all premises’ argument lists and goals from the corpus. The results of our method presented in Table 2 are obtained by first training the retrieval model and selecting the top-$100$ results as hard negative candidates. We then train the re-ranking model and use it to reorder the top-$20$ results from the retrieval stage. The experiments are conducted on 8 RTX 4080 servers.
Baselines As mentioned above, ReProver (Yang et al., 2024) is a model specifically trained on the Lean dataset, using formal states and premises. Therefore, it serves as our primary baseline. Moreover, we retrain the model using our dataset and following the setting in (Yang et al., 2024). In addition to ReProver, we compare against several other commonly used retrieval models that have demonstrated strong performance in their specific domains.
  • UniXcoder-base (Guo et al., 2022): It is one of the state-of-the-art code embedding models to transform code snippets into hidden vectors. It leverages multimodal data, such as code comments and abstract syntax trees (ASTs), to pre-train code representations.
  • E5-large-v2 (Wang et al., 2022): It is trained in a contrastive manner using weak supervision signals derived from a curated, large-scale text-pair dataset (referred to as CCPairs) and demonstrates strong performance on several English retrieval tasks.
  • BGE-m3 (Chen et al., 2024): It is distinguished for its versatility and excels in multiple functionalities, multilingual capabilities, and fine granularity retrieval tasks.
We fine-tune the three aforementioned baselines using our dataset. Following the parameters provided in (Chen et al., 2024), we set the learning rate to 1e-5 and use a linear scheduler. Due to server constraints, the batch sizes for the three models are set to 16, 8, and 6, respectively.
Metrics To evaluate the performance of various retrieval models, we utilize four widely adopted metrics: Precision, Recall, F1-Score, and nDCG. Since the nDCG considers the retrieved results’ relevance and position, it allows for multiple relevance levels. Following the setup in (Gao et al., 2024), we define the relevance criteria in Table 1, which provides the relevance scores for calculating nDCG.
@@
[…]
Match 1 Match to one of the premises used in the tactic.
@@
Data Split To evaluate the performance of the model across different feature datasets, we employ four distinct data split strategies. Here, we represent a proof as $P=\{t_{i}\}_{i=1}^{|P|}$, which is a sequence of tactics, and apply Proof Length and Premise Frequency to characterize the proof, i.e.,
@@
Each dataset contains 65,567 theorems for training, while the validation and test sets each contain 2,000 theorems.
@@
Table 3: Ablation studies on pre-training, tokenizer, and similarity calculation on Random split dataset for the CFR module. Here P means the model is pre-trained or not, T means the model uses a new tokenizer or not, S means similarity calculation.
@@
P T S R@1 R@5 R@10 P@1 P@5 P@10 F@1 F@5 F@10 n@1 n@5 n@10
$\times$ $\times$ Eq. (3) 4.63 20.09 29.05 8.35 7.45 5.62 5.96 10.87 9.42 0.2443 0.3503 0.4030
$\times$ $\checkmark$ Eq. (3) 4.93 20.49 30.33 8.04 7.50 5.79 6.11 10.98 9.72 0.2368 0.3493 0.4034
$\checkmark$ $\times$ Eq. (3) 6.31 23.81 34.21 10.34 8.77 6.59 7.84 12.82 11.05 0.2484 0.3655 0.4240
$\checkmark$ $\checkmark$ Eq. (2) 5.65 25.87 36.78 9.36 9.55 7.12 7.05 13.95 11.92 0.2382 0.3662 0.4238
$\checkmark$ $\checkmark$ Eq. (3) 5.94 25.81 37.70 9.80 9.53 7.24 7.40 13.92 12.15 0.2422 0.3674 0.4293
4.2 Main Results
After training different models on the training dataset and evaluating them on the corresponding test sets, we obtain the results presented in Table 2. From this table, we can observe that our method demonstrates exceptional performance, outperforming other baseline models. Specifically, UniXcoder-base (Guo et al., 2022), E5-large-v2 (Izacard et al., 2022), and BGE-m3 (Chen et al., 2024) exhibit generally poor performance, which can be attributed to the substantial discrepancy between the pre-training corpora used by these models (based on natural language data) and the Lean language. Although these models are fine-tuned, they cannot fully capture the core aspects of the Lean language due to Lean’s unique syntax and semantic features and thus suffer suboptimal performance. In contrast, the ReProver (Yang et al., 2024) model performs relatively well, likely due to its specifically designed parameter-tuning strategy. Furthermore, ReProver utilizes ByT5 (Xue et al., 2022), a byte-level tokenizer, which may allow the model to better handle the special symbols and structures in the Lean language. Our method, on the other hand, improves performance by retraining the tokenizer on the Lean corpus and pre-training the retrieval model, allowing it to better adapt to the syntax and semantic features of Lean.
@@ 4.2 Main Results
4.3 Ablation Studies
@@
Figure 4: The Recall@k of our method with or without re-ranking on each split, compared with ReProver.
Impacts of Pre-Training, Tokenizer, and Similarity Calculation We evaluate the impacts of pre-training the model, training the specific tokenizer for Lean, and the proposed similarity calculation in Table 3, respectively. To ensure a fair comparison, we filter out the impact of the re-ranking module and only compare the results provided by the retrieval module. This table shows a performance improvement when the model is pre-trained on the Lean corpus compared to not performing pre-training. This indicates that the pre-training process helps the model acquire semantic information about Lean in advance. The experiment also shows that if we only pre-train the model without training a new tokenizer, the performance when $k=1$ is slightly better but degrades a lot when $k=5$ or $10$. In addition, we evaluate the impacts of different similarity calculation methods on the retrieval results. The results in Table 3 emphasize the benefits of assessing similarity in a fine-grained manner.
@@
The Effect of Re-ranking We validate the effectiveness of using the proposed re-ranking module. As shown in Figure 4, when only the CFR module is used, our model outperforms the ReProver only for larger values of $k$ in Recall@k (e.g., $k\geq 10$). Applying the CAR module improves the recall of our model when $k<10$, so that our model can consistently outperform ReProver and users are likely to find useful premises in a relatively short list. Figure 5 presents a comparison for our model and other baselines in terms of model size and inference GFLOPs. Without re-ranking, our model achieves the lowest GFLOPs and it outperforms in Recall@1 most of the baselines except ReProver. When re-ranking the top-$5$ retrieval results by our CAR module, our model outperforms all the baselines and its GFLOPs and model size are comparable to those of ReProver.
Figure 5: The performance and efficiency comparison on Random split datasets. The size of each point reflects the parameter size of each model.
4.4 Robustness Experiments